# Skyrmion motion in a synthetic antiferromagnet driven by asymmetric spin wave emission

### v1

@@ IV.4 The effect of antiferromagnetic interlayer exchange coupling
With our new understanding of the coupling between skyrmion velocity and the frequency of the out-of-phase breathing mode, we now move to studying it as a function of the antiferromagnetic interlayer exchange coupling strength. The results of measuring the velocity as a function of frequency at different coupling strengths $\sigma_{\mathrm{RKKY}}$ are presented in Fig. 6(a). The data presented here has been trimmed around the frequency of the out-of-phase breathing mode for clarity, so the broad peak in velocity associated with the in-plane field is not shown. This however follows the same behaviour—albeit shifted in frequency—as the one for $\sigma_{\mathrm{RKKY}}=0.3\ \mathrm{mJ/m^{2}}$ shown in Fig. 4(a) and discussed in supplementary note LABEL:supp:sec:high_freq_mode. The extracted values of skyrmion velocity as a function of frequency are fitted to an asymmetric Lorentzian lineshape, chosen for its empirical match to the data, which allows us to extract a value for the peak velocity. This peak velocity as a function of $\sigma_{\mathrm{RKKY}}$ is then shown in Fig. 6(b).
@@
First, we define the skyrmion annihilation regime for low values of $\sigma_{\mathrm{RKKY}}$ below $\sim 0.03\ \mathrm{mJ/m^{2}}$. Here the skyrmions are stable in the static case and skyrmion motion is still possible when driving with AC magnetic fields. However, when it is driven close to the resonant frequency, the oscillations of the skyrmions at this field strength are sufficiently large to drive the annihilation of the skyrmions in at least one of the layers. Clearly the strength of the antiferromagnetic coupling is insufficient to stabilise the two skyrmions as a topological object.
@@
[[[[CoFeB/Ru/Pt/ Domain Walls -0.018 [30]
@@
V Electric Field Driven Skyrmion Motion
@@
We also drive the motion of the skyrmions with an electric field of the same magnitude and varying frequency. The measured velocities as a function of the electric field frequency are also plotted on the right hand axis of Fig. 7(a). Here we observe a strong peak in velocity that is well coupled to the lower frequency breathing mode. Unlike the magnetic field driven case, there is no observable motion at the higher frequency mode, but there is some motion around the frequency of the mode excited by the in-plane field (see supplementary note LABEL:supp:sec:high_freq_mode). To understand this we present in Figs. 7(b) and (c) simulations of the radius variation of the skyrmions when driven by a sinusoidal electric field with no in-plane magnetic field at each of the principal frequencies from Fig. 7(a). At the lower 14.3 GHz peak, the radii of the two skyrmions varies massively with this electric field strength of $\langle\Delta r_{\mathrm{Sk_{1}}}\rangle=\langle\Delta r_{\mathrm{Sk_{2}}}\rangle=13.6\pm 0.2$ nm. Our hypothesis in Sec. IV.3 was that to stimulate skyrmion motion, a minimum variation in the skyrmion radius—of 0.44 nm for these micromagnetic energy terms—is required. The variation in radius at the in-phase resonant frequency is clearly large enough to drive motion. Likewise, this also explains the absence of motion at the other resonant breathing mode. The average variation in the radii of the two skyrmions when driven by an electric field at 57.4 GHz are $\langle\Delta r_{\mathrm{Sk_{1}}}\rangle=0.1768\pm 0.0001$ nm and $\langle\Delta r_{\mathrm{Sk_{2}}}\rangle=0.1821\pm 0.0001$ nm respectively, well below the threshold for motion.
VI Conclusion
@@
Finally, we have compared our results to the case when driven by a microwave electric field. By doing this we have revealed both the true origin of the relationship between velocity and exchange coupling, and the reason for the lower frequency of skyrmion motion when driven by electric fields compared to magnetic fields.

### v2

@@ IV.4 The effect of antiferromagnetic interlayer exchange coupling
With our new understanding of the coupling between skyrmion velocity and the frequency of the out-of-phase breathing mode, we now move to studying it as a function of the antiferromagnetic interlayer exchange coupling strength. The results of measuring the velocity as a function of frequency at different coupling strengths $\sigma_{\mathrm{RKKY}}$ are presented in Fig. 6(a). The data presented here has been trimmed around the frequency of the out-of-phase breathing mode for clarity, so the broad peak in velocity associated with the in-plane field is not shown. This however follows the same behaviour—albeit shifted in frequency—as the one for $\sigma_{\mathrm{RKKY}}=0.3\ \mathrm{mJ/m^{2}}$ shown in Fig. 4(a) and discussed in supplementary note S2. The extracted values of skyrmion velocity as a function of frequency are fitted to an asymmetric Lorentzian lineshape, chosen for its empirical match to the data, which allows us to extract a value for the peak velocity. This peak velocity as a function of $\sigma_{\mathrm{RKKY}}$ is then shown in Fig. 6(b).
@@
First, we define the skyrmion annihilation regime for low values of $\sigma_{\mathrm{RKKY}}$ below $\sim 0.03\ \mathrm{mJ/m^{2}}$. Here the skyrmions are stable in the static case and skyrmion motion is still possible when driving with microwave magnetic fields. However, when it is driven close to the resonant frequency, the oscillations of the skyrmions at this field strength are sufficiently large to drive the annihilation of the skyrmions in at least one of the layers. Clearly the strength of the antiferromagnetic coupling is insufficient to stabilise the two skyrmions as a topological object.
@@
[[[CoFeB/Ru/Pt/ Domain Walls -0.018 [30]
@@
V Skyrmion Motion in Synthetic Ferrimagnets
[…]
In experimental SAFs, it is often difficult to grow magnetic layers with perfect compensation, especially when using layers of multiple compensations [30, 52]. To extend our understanding of the effects presented in this paper, we consider synthetic ferrimagnets, where we introduce a decompensation of the magnetisation of one of the two layers. The measurement of the breathing modes has already been demonstrated in Ref. [39], where simulation of a synthetic ferrimagnet (SFiM) was achieved by varying the $M_{\mathrm{s}}$ of one layer. Similarly, the current driven dynamics of skyrmions in angularly decompensated SAFs have been studied by varying the ration of $M_{s1}/M_{s2}$ of the two layers [53]. Here, we follow this method, and examine our system at two characteristic antiferromagnetic interlayer exchange coupling strengths, while varying the $M_{\mathrm{s}}$ of each layer individually by $\pm$30of its original value. We note that in most experimental SAFs the decompensation between the magnetisations of each layer is $\leq$10[30, 52]. Thus, our chosen range is adequate to describe experiments well, while also extending our understanding to higher decompensations. Throughout we conduct every simulation by fixing the $M_{\mathrm{s}}$ of one layer to its original value, and vary the $M_{\mathrm{s}}$ of the second layer. We use the labels $M_{s1}$ and $M_{s2}$ to denote these two values, and for every simulation we consider the case of varying each of these two quantities.
The breathing mode frequencies for an antiferromagnetic interlayer exchange coupling strength of $\sigma_{\mathrm{RKKY}}=-0.1\ \mathrm{mJ/m^{2}}$ and a variation of $M_{s1}$ by $\pm$30are shown in Figure 7(a) as a heatmap. The PSD here is taken for only layer 1 of the system, in order to be able to observe the signal of the in-phase breathing mode, however the behaviour of the mode frequencies is the same for the PSD of layer 2 and the whole system. In general, the resonant frequencies of both the in-phase and the out-of-phase modes both decrease as the value of $M_{s1}$ increases. This matches the results of Ref. [39], where it was shown that in compensated SAFs, the frequencies of the resonant modes decrease with $M_{\mathrm{s}}$. Likewise, our results match those considering SFiM systems [39], where it was shown that the frequencies of the resonant modes continue to follow this trend even when the variation in $M_{\mathrm{s}}$ is restricted to only one of the magnetic layers.
With this understanding of the behaviour of the two resonant modes, we move on to consider the velocity of the skyrmions in uncompensated systems. The velocity as a function of frequency for several representative values of $M_{s1}$ is shown in Figure 7(b). This data is for an antiferromagnetic interlayer exchange coupling strength of $\sigma_{\mathrm{RKKY}}=-0.1\ \mathrm{mJ/m^{2}}$. Here, each dataset is deliberately offset along the $y$-axis to be able to see the individual trends. It is clear that the velocity at the out-of-phase peak is lower than at full compensation (see Figure 6 for the measurement at this value of $\sigma_{\mathrm{RKKY}}$), and this decreases as a function of the decompensation of $M_{s1}$. Conversely, at the highest decompensations, there is a clear evidence of a small peak in velocity near the in-phase breathing mode frequency. These two effects are better demonstrated in Figure 7(c), which shows the measured velocity at each of these two resonant frequencies as a function of decompensation for both $M_{s1}$ and $M_{s2}$. There is a clear peak in velocity at the out-of-phase resonant frequency when $M_{s1}=M_{s2}$, with skyrmion velocity trailing off at decompensation either side of this value. For $M_{s1,2}$ compensations below 90%, there is no velocity data. This is due to the small value of $M_{\mathrm{s}}$. For these low $M_{\mathrm{s}}$ values, the increased field amplitude used to drive motion (compared to the excitation of the breathing modes in Figure 7(a)), annihilates the skyrmions.
As expected, the behaviour is reversed when considering the velocity at the in-phase resonant frequency. The velocity is at a minimum value when $M_{s1}=M_{s2}$, and increases away from this. This can be simply understood considering the relationship between skyrmion velocity and breathing amplitude from Section IV.3. As the difference between the $M_{\mathrm{s}}$ of the layers increases, so too will the difference in the oscillation amplitudes of each skyrmion. As we have discussed this is the main factor determining the skyrmion velocity, and so naturally the velocity increases at this resonant frequency with the increase in unbalance.
Finally, we reproduce this work for a second antiferromagnetic interlayer exchange coupling strength of $\sigma_{\mathrm{RKKY}}=-0.3\ \mathrm{mJ/m^{2}}$. The measurement of the breathing modes and the individual velocity as a function of frequency plots for each decompensation, analogous to Figures 7(a) and 7(b) are not shown here, but are instead summarised in supplementary note S5. The velocity of each of the resonant modes as a function of the decompensation is shown in Figure 7(d). As before, the skyrmion is annihilated when the $M_{\mathrm{s}}$ is below 95 % of the original value, demonstrating that this is an intrinsic effect of the low magnetisation, rather than a consequence of the weaker coupling present in Figure 7(c). We see similar behaviour at compensations near to $M_{s1}=M_{s2}$, with a peak in velocity at the out-of-phase resonant frequency at full compensation, and increasing velocity at the in-phase resonant frequency with increasing decompensation. However, for values of $M_{s1,2}\geq 120\%$, there is a large degree of variation in the velocity at the out-of-phase resonant frequency. This is due to hybridisation between the skyrmion resonant frequency and the spin wave mode of the canted spins discussed in Section IV.2 and Supplementary Note S2. Here the resonant frequency of the spin wave mode arising from the canted spins decreases faster with increasing $M_{s1,2}$ than the skyrmion breathing mode frequency. Because the two modes are closer in frequency at this interlayer exchange coupling strength, at the higher decompensations, the spin wave mode overlaps with the skyrmion’s out-of-phase resonant frequency, and the two effects interact to create more unpredictable spin wave emission. There is also a trace of this effect in Figures 7(b,c) for the lower interlayer coupling strength, however it only occurs at the very highest decompensation, and so is less noticeable.
VI Electric Field Driven Skyrmion Motion
@@
We also drive the motion of the skyrmions with an electric field of the same magnitude and varying frequency. The measured velocities as a function of the electric field frequency are also plotted on the right hand axis of Fig. 8(a). Here we observe a strong peak in velocity that is well coupled to the lower frequency breathing mode. Unlike the magnetic field driven case, there is no observable motion at the higher frequency mode, but there is some motion around the frequency of the mode excited by the in-plane field (see supplementary note S2). To understand this we present in Figs. 8(b) and (c) simulations of the radius variation of the skyrmions when driven by a sinusoidal electric field with no in-plane magnetic field at each of the principal frequencies from Fig. 8(a). At the lower 14.3 GHz peak, the radii of the two skyrmions varies massively with this electric field strength of $\langle\Delta r_{\mathrm{Sk_{1}}}\rangle=\langle\Delta r_{\mathrm{Sk_{2}}}\rangle=13.6\pm 0.2$ nm. Our hypothesis in Sec. IV.3 was that to stimulate skyrmion motion, a minimum variation in the skyrmion radius—of 0.44 nm for these micromagnetic energy terms—is required. The variation in radius at the in-phase resonant frequency is clearly large enough to drive motion. Likewise, this also explains the absence of motion at the other resonant breathing mode. The average variation in the radii of the two skyrmions when driven by an electric field at 57.4 GHz are $\langle\Delta r_{\mathrm{Sk_{1}}}\rangle=0.1768\pm 0.0001$ nm and $\langle\Delta r_{\mathrm{Sk_{2}}}\rangle=0.1821\pm 0.0001$ nm respectively, well below the threshold for motion. This is a marked difference between the case of driving with an electric field and a magnetic field. We suggest that this arises from the difference in the mechanism of excitation. In the case of the electric field, this is simulated through a change in the anisotropy, which affects the radius of both skyrmions equally. On the other hand, when the breathing modes are excited with a magnetic field, the Zeeman coupling will have an opposite effect on each of the skyrmions due to the antiferromagnetic coupling. Thus, in the magnetic field case, at the in-phase resonant mode, the Zeeman effect will always suppress the oscillation of one of the two skyrmions and so the overall amplitude is smaller. There is no such competition in the electric field/anisotropy case, and so at this resonant mode maximum oscillations are allowed.
VII Skyrmion Motion Driven By Perpendicular Current Densities
Figure 9: (a). Illustration of the current perpendicular to plane (CPP) spin transfer torque (STT) geometry, with the direction of the current illustrated with the arrow labeled $J_{c}$ and the polarisation of the fixed layer illustrated with $\mathbf{m_{p}}$. (b). Spectral response of skyrmion system when stimulated with a CPP-STT with a current density of $1\times 10^{9}\mathrm{A/m^{2}}$. (c). Skyrmion velocity when driven at the OOP breathing mode resonant frequeny of 57.6 GHz found from (b) as a function of current density. (d). Sweeps of the skyrmion velocity as a function of frequency for several representative current densities chosen from (c).
In this section, we study the effects of the skyrmion breathing mode excitation by an electric current via a spin-torque mechanism (instead of a microwave magnetic field) in a fully compensated SAF. It has already been shown that a Slonczewski-type STT [55, 56, 57] can be used to qualitatively reproduce the effects of the microwave magnetic field when stimulating breathing modes [39]. Here, we extend this to considering the effects of an STT on the skyrmion velocity in the presence of a static in-plane magnetic field.
To model a Slonczewski type STT, an additional torque term is added to the right hand side of the LLG (Equation 1), given by:
$\vec{\mathbf{\tau}}_{SL}=\beta\dfrac{\epsilon-\alpha\epsilon^{\prime}}{1+\alpha^{2}}\Big{[}\vec{\mathbf{m}}\times\Big{(}\vec{\mathbf{m}}_{P}\times\vec{\mathbf{m}}\Big{)}\Big{]}-\dfrac{\epsilon^{\prime}-\alpha\epsilon}{1+\alpha^{2}}\Big{(}\vec{\mathbf{m}}\times\vec{\mathbf{m}}_{P}\Big{)}.$ (11)
Here, $\vec{\mathbf{m}_{P}}$ is the magnetisation of the fixed layer—which is not explicitly simulated—and $\epsilon$ and $\epsilon^{\prime}$ represent the effective spin polarisation terms for the damping-like and field-like torques, respectively. We can write $\epsilon$ as
$\epsilon=\dfrac{P\Lambda^{2}}{(\Lambda^{2}+1)+(\Lambda^{2}-1)(\vec{\mathbf{m}}\cdot\vec{\mathbf{m}_{P}})},$ (12)
where P is the spin polarisation and the Slonczewski $\Lambda$ parameter is a dimensionless parameter of the model. Finally, the last dimensionless parameter $\beta$ is given by
$\beta=\dfrac{\hbar J_{c}}{M_{\mathrm{s}}ed},$ (13)
where $J_{c}$ is the current density and $d$ is the film thickness. For all subsequent results, we simulate a skyrmion in a system with an antiferromagnetic interlayer exchange coupling energy of $\sigma_{\mathrm{RKKY}}=-3\times 10^{-4}\ \mathrm{J/m^{2}}$, and driven by a pure damping-like torque ($\epsilon\neq 0,\epsilon^{\prime}=0$). For a Slonczewski-type STT the current density $J_{c}$ is applied along the out-of-plane axis, in what is described as current perpendicular to plane (CPP). The direction of the fixed layer is also set along the $+z$ direction. A schematic of the simulated system along with the direction of the current density and the polarisation of the fixed layer is shown in Figure 9(a). For all simulations the STT is considered to be present in both ferromagnetic layers. Considering the spin diffusion length $l_{\mathrm{Ru}}$ of ruthenium, the material typically chosen as the nonmagnetic spacer layer, of 14 nm [58], this set-up is valid [39].
Figure 9(b) shows the spectral response of the system when excited with a sinc pulse of current density of $1\times 10^{9}\ \mathrm{A/m^{2}}$. The frequencies of both the in-phase and out-of-phase modes reproduce those shown in Figure 2(a). As mentioned, the excitation of skyrmion breathing modes was studied extensively in Ref. [39], where they showed that for their chosen magnetic parameters the breathing modes occurred at the same frequencies regardless of whether they were excited by a magnetic field or a current density. Furthermore, they demonstrated that this was independent of the nature of the STT; demonstrating the same effects with field-like STTs ($\epsilon=0,\epsilon^{\prime}\neq 0$) or a mixture of damping-like and field-like STTs ($\epsilon\neq 0,\epsilon^{\prime}\neq 0$).
Figure 9(c) shows the velocity of the skyrmion as a function of the current density when driven by an oscillating current density with a frequency of 57.6 GHz—which was obtained by fitting to the peak of the out-of-phase breathing mode in Figure 9(b). The current density was varied in a quasi-logarithmic fashion to allow a sampling of a wide range of magnitudes. At current densities much above $5\times 10^{11}\ \mathrm{A/m^{2}}$, significant deformation of the skyrmion was observed, before eventually breaking down at current densities in the $10^{12}\ \mathrm{A/m^{2}}$ (not shown). We extend this to examine the effects of the skyrmion velocity as a function of frequency at current densities of $5\times 10^{10}$, $1\times 10^{11}$ and $2\times 10^{11}\ \mathrm{A/m^{2}}$ in Figure 9(d). The shape of the $v(f)$ is broadly similar to that presented in Figure 4. However, some important differences should be noted. In a similar way to the magnetic-field driven case, we observe little-to-no evidence of any motion at the resonant frequency of the in-phase breathing mode. This is because the CPP-STT is acting as an effective magnetic field on the skyrmions, and so we expect any behaviour to be consistent with the case of driving with a simple magnetic field. Likewise, the relative velocities of the skyrmions driven at the resonant frequency of the out-of-phase breathing mode and at the resonant frequency of the spin wave mode arising from the canted spins are different. While the absolute magnitude of the velocity depends on the strength of the current density or magnetic field, this relative difference is of note. It arises because a current density along the $z$-axis, with a spin polarisation also along the $z$-axis will induce an effective field that acts on the in plane spins. In the case of the skyrmions, this are only the spins in the narrow band around the skyrmion where the magnetisation rotates from up to down (and vice-versa for the skyrmion in the other layer). When the spin wave emission is arising from the hybrid skyrmion breathing and resonance of the canted spins of the system, then we would expect this to be coupled much better to the STT, and so increase the velocity of the skyrmion, as we see for each of the current densities in Figure 9(d).
VIII Conclusion
@@
We have also extended our study to consider synthetic antiferromagnets, which is relevant to experimental studies where a perfect compensation of moments is difficult to achieve. We show that as expected, the velocity is at a maximum when the system is fully compensated, however motion is possible over a wide range of decompensation, and this is true for low and relatively high coupling strength (in comparison to what has been reported experimentally).
Finally, we have compared our results to the case when driven by a microwave electric field and oscillating current densities. By doing this we have revealed both the true origin of the relationship between velocity and exchange coupling, and the reason for the lower frequency of skyrmion motion when driven by electric fields compared to magnetic fields, as well as demonstrating the wide range of mechanisms through which this spin wave emission skyrmion motion can be driven.